In 2021, Junica Co. acquired land costing ore for a total cost of ₱4,000,000 Exploration costs amounted to ₱6,000,000 while intangible development cost amounted to ₱5,000,000. Movable tangible equipment costs amounted to ₱1,000,000. Geologist estimated that the total units estimated to be extracted is 4,000,000. It is estimated that 400,000 units will be extracted to be year during the useful life of the wasting assets.

The movable equipment has a useful life of 10 years while the immovable equipment has an estimated useful life of 5 years.

Actual units extracted during 2021 are 500,000 units.

Questions:

Based on the above data, answer the following:

1. How much is the depletion in 2021?

2. How much is the 2021 depreciation of the movable equipment?

3. How much is the 2021 depreciation of the immovable equipment?



1
Expert's answer
2021-06-17T09:32:47-0400

Solution:

a.). Depletion in 2021:


Depletion = "\frac{Total \;costs - Salvage \;value}{Estimated \;no. \;of\; units}\times No.\; of \;units\; extracted"


Total costs = 4,000.000 + 6,000,000 + 5,000,000 = 15,000,000

Estimated no. of units = 4,000,000 units

No. of units extracted in 2021 = 500,000 units


Depletion in 2021 = "\frac{15,000,000 - 0}{4,000,000}\times 500,000 = 1,875,000"


Depletion in 2021 = ₱1,875,000


b.). 2021 depreciation of the movable equipment:


Depreciation = "\frac{Original \;cost - Salvage \;value}{Estimated \;production \; capability}\times No.\; of \;units\; produced \; for\; the\; year"


Original cost of the movable equipment = 1,000,000

Estimated production capability = 400,000 "\times"10 = 4,000,000 units

No. of units produced in 2021 = 500,000 units


Depreciation = "\frac{1,000,000 - 0}{4,000,000}\times 500,000 = 125,000"


Depreciation = ₱125,000


c.). b.). 2022 depreciation of the immovable equipment:


Depreciation = "\frac{Original \;cost - Salvage \;value}{Estimated \;production \; capability}\times No.\; of \;units\; produced \; for\; the\; year"


Original cost of the immovable equipment = 3,000,000

Estimated production capability = 400,000 "\times"15 = 6,000,000 units

No. of units produced in 2021 = 500,000 units


Depreciation = "\frac{3,000,000 - 0}{6,000,000}\times 500,000 = 250,000"


Depreciation = ₱250,000